OpenStudy (lexi_loves):

I need to find the slope of y-2=-3(x+1)

OpenStudy (lexi_loves):

Can some one help me?!

pooja195 (pooja195):

Let's make it into slope intercept form y-2=-3(x+1) Start by distribtuing -3 to the x and 1 what do you get?

OpenStudy (lexi_loves):

@Abstract_Undercover @aduller @just_one_last_goodbye

pooja195 (pooja195):

.-.

OpenStudy (lexi_loves):

you get y-2=-3x+1

Nnesha (nnesha):

\[y-y_1=m(x-x_1)\] is a point slope form where m is the slope

pooja195 (pooja195):

Good you dont even need to add the 2 y=mx+b m=slope what number is attatched to the x?

pooja195 (pooja195):

*attached

OpenStudy (lexi_loves):

-3

pooja195 (pooja195):

Good that's out slope.

pooja195 (pooja195):

*our

OpenStudy (lexi_loves):

ohh ok so how would i graph that?

pooja195 (pooja195):

Graph what?

OpenStudy (lexi_loves):

how do i graph the slope?

pooja195 (pooja195):

Slope=rise/run

pooja195 (pooja195):

-3/1

pooja195 (pooja195):

You need a y intercept you get y-2=-3x+1 Add 2 to both sides y=-3x+2 Now you can properly graph it

OpenStudy (lexi_loves):

ohh ok i see!!! thank you so much!! :)

pooja195 (pooja195):

yw =D